OpenStudy (anonymous):

Find the coordinates of the other endpoint if the midpoint is M(8, 2) and the other endpoint is P(5, 6).

OpenStudy (ivettef365):

use the midpoint formula x1 + x2 y1 + y2 ------- , -------- 2 2

OpenStudy (akashdeepdeb):

Let the other one be (x,y) So the mid point is given by mid-point formula (8,2) = [(x+5)/2,(y+6)/2] So you get 2 equations! 8 = (x+5)/2 And 2 = (y+6)/2 Solve them and get the other point!! :D Getting me? :)

OpenStudy (anonymous):

thank you! that's what I thought were were supposed to do, but I couldn't remember the formula

OpenStudy (akashdeepdeb):

:)